Let a=cos(π/3), b=cos(1radian), c=cos(1degree), we have that


A.c a


1
Expert's answer
2021-08-31T03:42:25-0400

The function f(t)=cos(t) is decreasing if 0"\le t \le \pi" .

We have 0<t1=1 degree <t2=1 radian "\approx 57^\circ <\pi/3=60^\circ" , therefore

c=f(t1)>b=f(t2)>a=f(t3). So we have B.a<b<c.